# CCDM: Continuous Conditional Diffusion Models for Image Generation

### v2

@@ IV-A Experimental Setup
For the $192\times 192$ experiment, we focus exclusively on comparing the four CCGM models, including CcGAN (SVDL+ILI), Dual-NDA, CcDPM, and CCDM.
@@
Implementation Setup. In our implementation of class-conditional models, we bin the regression labels of four datasets into various classes: 150 classes for RC-49, 60 classes for UTKFace, 221 classes for Steering Angle, and 100 classes for Cell-200. To facilitate comparisons, we leveraged pre-trained models provided by Ding et al.[ding2024turning] for experiments on UTKFace ($64\times 64$ and $128\times 128$) and Steering Angle ($64\times 64$ and $128\times 128$) datasets, rather than re-implementing all candidate methods (excluding CCDM). Additionally, Ding et al.[ding2023ccgan] offered pre-trained ReACGAN for RC-49 and CcGAN (SVDL+ILI) for UTKFace ($192\times 192$). For other candidate models, we re-implemented them for comparison. When implementing CcGAN (SVDL+ILI) and Dual-NDA on Cell-200, we used DCGAN [radford2015unsupervised] as the backbone and trained with the vanilla cGAN loss [mirza2014conditional]. For RC-49, we employed SAGAN [zhang2019self] as the backbone and trained with hinge loss [lim2017geometric] and DiffAugment [zhao2020differentiable]. For CCDM, we set $p_{\text{drop}}=0.1$ in Remark 3 and $\gamma=1.5$ in Eq. (21) for most experiments, except for UTKFace ($128\times 128$ and $192\times 192$), where we increased $\gamma$ to 2.0 for better label consistency. The total time step $T$ is fixed at 1000 during training, while the sampling time step $T^{\prime}$ varies: 250 for the $64\times 64$ experiments, and 150 for all other experiments. The hyperparameters $\kappa$ and $\sigma_{\delta}^{2}$ were set according to the rule of thumb in Remark 3 of [ding2023ccgan]. The training configurations of CcDPM were largely similar to those of CCDM, with the exception that CcDPM utilized $\bm{\epsilon}$-prediction loss and employed the soft vicinal weight. The one-step sampler, DMD2-M, is implemented only in the RC-49 ($64\times 64$) and Steering Angle ($64\times 64$) experiments, where it is compared with the DDPM and DDIM samplers. Detailed implementation setups are provided in Appendix LABEL:sec:supp_detailed_train_setups.
@@
Evaluation Setup. Consistent with the work of Ding et al. [ding2021ccgan, ding2023ccgan, ding2023efficient, ding2024turning], we predefine $m_{c}$ distinct points within the regression label range for each experiment, which we refer to as evaluation centers. For each evaluation center, we generate $N_{c}^{g}$ fake images, resulting in a total of 179,800, 60,000, 100,000, and 200,000 images generated by each candidate method in the RC-49, UTKFace, Steering Angle, and Cell-200 experiments, respectively. These images are evaluated using one overall metric and three separate metrics. The overall metric is the Sliding Fréchet Inception Distance (SFID) [ding2023ccgan], calculated as the average FID score across evaluation centers, with the standard deviation of these FIDs reported in Tabel  LABEL:tab:main_results. Additionally, we use Naturalness Image Quality Evaluator (NIQE) [mittal2012making] for visual fidelity, Diversity [ding2023ccgan] for image variety, and Label Score [ding2023ccgan] for label consistency. Note that the Diversity score cannot be computed for the Cell-200 dataset due to the lack of categorical image labels. For SFID, NIQE, and Label Score, lower values indicate better performance, while higher Diversity scores are preferred for greater variety in generated samples. For detailed evaluation setups, please refer to Appendix LABEL:sec:supp_detailed_test_setups.
@@ IV-B Experimental Results
We compare candidate methods across seven settings involving four datasets and three resolutions, as shown in Table LABEL:tab:main_results. The performance of DMD2-M against three other samplers is summarized in Table LABEL:tab:ab_effect_samplers. Visual results are presented in Fig. LABEL:fig:line_graphs_rc49 and Fig. LABEL:fig:example_fake_images in the Appendix. Our analysis of the experimental results highlights the following findings:

### v3

@@ IV-A Experimental Setup
  • A text-to-image diffusion model: We include Stable Diffusion v1.5 (SD v1.5) [rombach2022high] in our comparative evaluation on RC-49 and UTKFace, performing full fine-tuning of the official Hugging Face checkpoint (stable-diffusion-v1-5).
@@
For the $192\times 192$ experiment, we focus exclusively on SD v1.5, CcGAN (SVDL+ILI), Dual-NDA, CcDPM, and CCDM.
@@
Implementation Setup. In our implementation of class-conditional models, we bin the regression labels of four datasets into various classes: 150 classes for RC-49, 60 classes for UTKFace, 221 classes for Steering Angle, and 100 classes for Cell-200. To facilitate comparisons, we leveraged pre-trained models provided by Ding et al.[ding2024turning] for experiments on UTKFace ($64\times 64$ and $128\times 128$) and Steering Angle ($64\times 64$ and $128\times 128$) datasets, rather than re-implementing all candidate methods (excluding CCDM). Additionally, Ding et al.[ding2023ccgan] offered pre-trained ReACGAN for RC-49 and CcGAN (SVDL+ILI) for UTKFace ($192\times 192$). For other candidate models, we re-implemented them for comparison. When implementing CcGAN (SVDL+ILI) and Dual-NDA on Cell-200, we used DCGAN [radford2015unsupervised] as the backbone and trained with the vanilla cGAN loss [mirza2014conditional]. For RC-49, we employed SAGAN [zhang2019self] as the backbone and trained with hinge loss [lim2017geometric] and DiffAugment [zhao2020differentiable]. For CCDM, we set $p_{\text{drop}}=0.1$ in Remark 3 and $\gamma=1.5$ in Eq. (21) for most experiments, except for UTKFace ($128\times 128$ and $192\times 192$), where we increased $\gamma$ to 2.0 for better label consistency. The total time step $T$ is fixed at 1000 during training, while the sampling time step $T^{\prime}$ varies: 250 for the $64\times 64$ experiments, and 150 for all other experiments. The hyperparameters $\kappa$ and $\sigma_{\delta}^{2}$ were set according to the rule of thumb in Remark 3 of [ding2023ccgan]. The training configurations of CcDPM were largely similar to those of CCDM, with the exception that CcDPM utilized $\bm{\epsilon}$-prediction loss and employed the soft vicinal weight. The one-step sampler, DMD2-M, is implemented only in the RC-49 ($64\times 64$) and Steering Angle ($64\times 64$) experiments, where it is compared with the DDPM and DDIM samplers. Stable Diffusion v1.5 (an advanced text-to-image model in our comparison) is fine-tuned on both RC-49 and UTKFace datasets, using the prompt templates “a photo of a chair at yaw angle [ANGLE] degrees” and “a portrait of the face of [AGE] year old”, respectively. The detailed implementation configurations are available in our GitHub repository.
@@
Evaluation Setup. Consistent with the work of Ding et al. [ding2021ccgan, ding2023ccgan, ding2023efficient, ding2024turning], we predefine $m_{c}$ distinct points within the regression label range for each experiment, which we refer to as evaluation centers. For each evaluation center, we generate $N_{c}^{g}$ fake images, resulting in a total of 179,800, 60,000, 100,000, and 200,000 images generated by each candidate method in the RC-49, UTKFace, Steering Angle, and Cell-200 experiments, respectively. These images are evaluated using one overall metric and three separate metrics. The overall metric is the Sliding Fréchet Inception Distance (SFID) [ding2023ccgan], calculated as the average FID score across evaluation centers, with the standard deviation of these FIDs reported in Table  LABEL:tab:main_results. Additionally, we use Naturalness Image Quality Evaluator (NIQE) [mittal2012making] for visual fidelity, Diversity [ding2023ccgan] for image variety, and Label Score [ding2023ccgan] for label consistency. Note that the Diversity score cannot be computed for the Cell-200 dataset due to the lack of categorical image labels. For SFID, NIQE, and Label Score, lower values indicate better performance, while higher Diversity scores are preferred for greater variety in generated samples. For detailed evaluation setups, please refer to Appendix LABEL:sec:supp_detailed_test_setups.
@@ IV-B Experimental Results
We evaluate candidate methods across seven experimental settings spanning four datasets and three resolutions, with complete quantitative results documented in Table LABEL:tab:main_results. Representative visual results are shown in Fig. LABEL:fig:line_graphs_rc49 and Fig. LABEL:fig:example_fake_images. Sampling cost comparisons for CCGM methods are performed on RC-49, with quantitative results detailed in Table LABEL:tab:ab_sampling_cost. The performance of DMD2-M against three other samplers is summarized in Table LABEL:tab:ab_effect_samplers. Our analysis of the experimental results highlights the following findings:
@@
  • While CCDM ($T^{\prime}=250$) achieves the lowest SFID scores, this comes at the cost of substantially higher computational requirements. Although DMD2-M demonstrates a viable trade-off between image quality and sampling efficiency (reducing both time and memory usage), developing more efficient sampling algorithms for CCDM remains an important direction for future research.